January 29th 2021

From Hall A Wiki
Jump to: navigation, search
  • Updates
    • HCal
      • still taking cosmics
    • Bigbite hodoscope
      • working on software SBS-online, still working on root tree with SBS online. Basic decoder working
    • Bigbite shower
      • different trigger for BB shower - cosmics trigger coming from shower sum with scintillator - FADC only look backs in time - need to delay trigger currently using VME - will have to delay the trigger for the FADC if use same cables
      • FADC order in - expect get those 1 or 2 months
      • need integral , time and pedestal subtraction - mode 7 gives pedestal - time + integral + pedestal - mode 9 in Hall C ? need to check and could add to Hall B firmware
      • can use delay from TI - can delay FADC trigger with software 1 us scale
    • HRS
      • all 3 Fastbus crate working stand alone - there was a bug in TDC readout list for 1877 - need to do podd module to look at data - looking at FADC crate - will use as master with 3 Fastbus as slave - need to add the transceiver to connect the TI from Fastbus - will take RHRS FADC for LHRS -

6 FADCs - need to plan for aerogel -

      • working on Centos 7 install - will compare with the TEDF setup
      • detector not in yet in stack
    • GEM INFN : running pedestal smoothly - will take cosmics again next week - loose cable resolved on Monday for MPD17 -

added wire holder : at patch panel - short HDMI or long HDMI cable become loose - unplugging and replugging resolve the issue - can be replaced if needed - analog cable - reset LV at the same time - check if reset can fix it

    • UVA GEM setup
      • taking cosmics with new layers , connections were fixed and taking data now - one new MPD giving readout error in same crate as other were replaced - other previous 2 error were fixed with replacing MPD
      • computer xfs error : not issue since Thir fixed it with xfs_repair
      • Arunda starting to make inventory of MPDs - about 3 MPDs have issues - some APVs might be bad - inventory in process - when we have the lot of MPD send to Paolo for check and maybe send to company to be fixed - no plan for visit in US

, could meet with Paolo and electronics group to try to fix remotely

      • new MPD order , Nilanga in contact with company , NSF money available
      • Update GEM readout from Ben : bugs addressed with new firmware, running in TEDF right - issues appear when running buffering - corruption ( cannot read and write at same time ) - 800 Hz no zero suppression - 1.8 KHz with zero suppression ( expect 2.7 kHZ when optimized from simulation ) - Will setup the common mode supression and zero
      • 4.5 KHz with VTP readout

supression - will be ported to VTP - look at DAQ rate as function of occupancy - expect 30 % occupancy

      • event viewer  : to look at evio file by Xinzhan
    • CDet
      • new module on the bench being setup , PMTs installed next week and will start taking cosmics data
      • VETROC received
      • VXS crate received
      • clrlpc
      • need layout for CDet - will work with Peter
      • David set up CODA on clrlpc, will get back to it now hardware is available
  • Helicity : Paul might be able to help, new helicity board - recording the helicity pattern for each trigger - pseudorandom - board going out for layout soon will check with Ed for status - expect boards in spring - will compare with board and usual helicity readout for cross checks
  • GEn RP : setting up FADCs readout in ESB -

GEM testing for GEN RP - two groups on for Large Angle should move to ESB ( 1 to 2 months time scale ) - and can try to readout out with hodoscope - Rest of GEM will be assembled in the Hall- option to mount RP in cleanroom so GEM setup already set up ( will check ) - extra space for inline GEM - could be added to the cleanroom

  • Still need CPU in crate for CPUs
    • get new CPUs for the VME crate - will ask for quote and check with Thia